OpenStudy (anonymous):

Solve 3(2x - 8) = 12

OpenStudy (anonymous):

Multiply using distributive property

OpenStudy (anonymous):

x=6

OpenStudy (anonymous):

x=6

OpenStudy (anonymous):

Divide by 3 both sides, 2x - 8 = 4 2x = 12 x = 6 is the answer...

OpenStudy (anonymous):

a(b+c) = ab + ac